OpenStudy (anonymous):

In 1980 median family income was about$16,000, and in 2000 it was bout $40,000. (a) find the slope of the line passing through points 1980,16000 and (2000,40000). (b) Interpret the slope rate of change. (c) If this trend continues, estimate the median family income in 2006

OpenStudy (anonymous):

Do you know the formula for slope?

OpenStudy (anonymous):

let y2-y1 over x2-y2

OpenStudy (anonymous):

and divide by 3

OpenStudy (anonymous):

\[M=y _{2}-y_{1}/x_{2}-x_{1}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright, your two points are (1980, 16000) and (2000, 40000) whats the slope?

OpenStudy (anonymous):

the slope do i subtract 40,000 from 16,000

OpenStudy (anonymous):

and divide by 3

OpenStudy (anonymous):

No there is no dividing by 3, use the slope formula

OpenStudy (anonymous):

well you have to subtract the two top numbers and the instructions says divide by 3

OpenStudy (anonymous):

i am going to look again to see what two numbers for the slope

OpenStudy (anonymous):

You do 40000-16000/2000-1980

OpenStudy (anonymous):

ok i got to calculate

OpenStudy (anonymous):

24,000/20

OpenStudy (anonymous):

Which is 1200

OpenStudy (anonymous):

yes i got that too

OpenStudy (anonymous):

Alright do you know how to do point-slope form?

OpenStudy (anonymous):

i believe the family rate decreases by 12,000 between 1980-2000

OpenStudy (anonymous):

No, the median family income increased by 24000 between 1980 and 2000. When it says to interpret the slope that means to say what it means in the context of this problem. It means every year the median income increases by $1200

OpenStudy (anonymous):

oh ok i see so the family income of 12,000 will increase in 2006

OpenStudy (anonymous):

silly me thank you

OpenStudy (anonymous):

Wait, for part c we have to do something else

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

this is the formula we use to get an equation that we can use to estimate the income in 2006. \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

SO WEHAVE TO FIND THE FAMILY INCOME FOR 2006

OpenStudy (anonymous):

So which point do you want to use to plug into the formula

OpenStudy (anonymous):

SO THE M=12000 ON HERE?

OpenStudy (anonymous):

no 1,200. you added a 0

OpenStudy (anonymous):

Do you want to use the point (1980, 16000)?

OpenStudy (anonymous):

16,000-12,000

OpenStudy (anonymous):

when we plug everything in, itll look like y-16000=1200(x-1980)

OpenStudy (anonymous):

Y-16000=12000X-23760000 i think this is wrong but i multiplied12000 into whats in the brackets

OpenStudy (anonymous):

No thats what you do first, but you keep adding a 0. the slope is 1,200, not 12,000

OpenStudy (anonymous):

y-16000=1200x-237600

OpenStudy (anonymous):

You keep forgetting or adding zeros. y-16000=1200x-2376000

OpenStudy (anonymous):

ok well i think im not getting this second part

OpenStudy (anonymous):

You have the right idea you just keep adding or forgetting zeros. after we get that we subtract 16000 from both sides

OpenStudy (anonymous):

ok subtract 16000 and after i divide

OpenStudy (anonymous):

Im sorry, add 16000, not subtract. you end up with: y=1200x-2360000

OpenStudy (anonymous):

where x stands for the year and y stands for the median income. What would you plug in for x since you want to find the income for 2006?

OpenStudy (anonymous):

oh ok add 16000 because it is already a negative so change to add

OpenStudy (anonymous):

Yep so if you want to find the income for 2006, what do you plug in for x?

OpenStudy (anonymous):

y=1200(20)-2360000

OpenStudy (anonymous):

2000-1980=20

OpenStudy (anonymous):

x=20

OpenStudy (anonymous):

No, x represents the year, if you want to find the income for 2006, you plug in 2006 for x. y=1200(2006)-2360000

OpenStudy (anonymous):

ok i calculate now for answer

OpenStudy (anonymous):

47,200

OpenStudy (anonymous):

Thats correct

OpenStudy (anonymous):

omg this is hard

OpenStudy (anonymous):

Well now you're all done =D

OpenStudy (anonymous):

thanks i guess its hard with my infant aggravating me thank u